Is log((n^2)!) , ?(?log?)? Please show steps of proof. (use limits if possible)
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

By limit theorem

f(n)=w(g(n)) iff f(n)/g(n)=inf for n tending inf

f(n)=log(n^2!)

We can say that n^2!>=(n^2/2)^(n^2/2)=(n^2/2)^(n^2/2)

So,

f(n)=log(n^2!)>=log((n^2/2)^(n^2/2))=(n^2/2)*log(n^2/2)-------------------------eq1

g(n)=n*log(n)

So,

f(n)/g(n)=log(n^2!)/(n*log(n))

So, from equation 1

f(n)/g(n)>=(n^2)*log(n^2/2)/(2*n*log(n))=(n*log(n^2)-n*log(2))/(2*log(n))=(2*n*log(n)-n*log(2))/(2*log(n))

f/g>=n-n*log(2)/(2*log(n))

Since for n tending inf

f/g>=inf

which means f(n)/g(n)=inf

So, f(n)=w(g(n))

f(n)/g(n)=

Kindly revert for any queries

Thanks.